OpenStudy (anonymous):

Which fractions are greater than -5/12? A. -5/11 B. -3/8 C. -4/8 D.- 1/4 @iGreen @PaulaLovesSchool13 @damonte1 @iPurple please help Medal!!

OpenStudy (anonymous):

Choose exactly 2 answers that are correct.

OpenStudy (igreen):

-5 / 12 = -0.42 -5 / 11 = -0.45 -3 / 8 = -0.375 -4 / 8 = -0.5 -1 / 4 = -0.25 Remember, with negative numbers, the BIGGER the negative number seems, it's actually SMALLER.

OpenStudy (igreen):

So what's are your answers gonna be? @MeganNagem2002

OpenStudy (anonymous):

I'm going to go with a and b

OpenStudy (igreen):

No.. A is wrong.

OpenStudy (igreen):

-0.45 looks like it's bigger than -0.42, so it's actually smaller.

OpenStudy (igreen):

B is correct.

OpenStudy (anonymous):

b and d?

OpenStudy (igreen):

Yep!

OpenStudy (anonymous):

thanks sooo much